The mean lifetime for a sample of 125 lamps is 1205 hours with standard deviation 105 hours. However, the company claims that their lamps average lifetime is difference from 1300 hours. Test the claim at 1% level of significance.


1
Expert's answer
2021-06-14T16:14:40-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=1300"

"H_1:\mu\not=1300"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha=0.01," and the critical value for a two-tailed test for "\alpha=0.01" and "df=n-1=125-1=124" degrees of freedom is "t_c=2.61606." The rejection region for this two-tailed test is "R=\{t: |t|>2.61606\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{1205-1300}{105/\sqrt{125}}\approx-10.11555"

Since it is observed that "|t|=10.11555>2.61606=t_c," it is then concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is different than 1300, at the "\alpha=0.01" significance level.


Using the P-value approach: The p-value for two-tailed, "t=-10.11555, df=124," "\alpha=0.01" is "p=0," and since "p=0<0.01," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than 1300, at the "\alpha=0.01" significance level.